OpenStudy (anonymous):

4x - 3y = 9 find the slope

OpenStudy (anonymous):

3/4

OpenStudy (anonymous):

crap. 4/3 lol

OpenStudy (nenadmatematika):

no, it's 4/3

OpenStudy (anonymous):

take derivative of y w.r.t x 4-3dy/dx=0 -3dy/dx=-4 dy/dx=4/3 is required slope